# The viscoelastic rheology of transient diffusion creep

### v1

@@ A.1 Elasticity
$\nabla\cdot\underline{\underline{\boldsymbol{\sigma}}}=\mathbf{0}$ (11)
@@
$\sigma_{ij}=\lambda e_{kk}\delta_{ij}+2\mu e_{ij}$ (12)
@@
$e_{ij}=\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right).$ (13)
@@ A.2 Grain boundary diffusion
On the boundaries between grains the elastic displacement is discontinuous. Diffusion of atoms and vacancies along the grain boundaries leads to plating out or removal of material. Grain boundary diffusion is described by [e.g. 7, section 2.2]
$\left[\dot{\mathbf{u}}\cdot{\mathbf{n}}\right]+\Omega\delta D^{\text{gb}}_{\text{v}}\nabla_{\perp}^{2}c=0$ (14)
@@
$c=c_{0}\left(1+\frac{\Omega p}{kT}\right)$ (15)
@@
$p\equiv{\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\sigma}}}\cdot{\mathbf{n}}\quad\text{on }S,$ (16)
where $S$ is the grain-boundary interface. (15) can be substituted into (14) to write
$\left[\dot{\mathbf{u}}\cdot{\mathbf{n}}\right]+\frac{\Omega\delta D^{\text{gb}}}{kT}\nabla_{\perp}^{2}p=0,\$ (17)
@@ A.3 Grain boundary sliding
${\mathbf{n}}\times\underline{\underline{\boldsymbol{\sigma}}}\cdot{\mathbf{n}}=\frac{\eta_{\text{gb}}}{\delta}[{\mathbf{n}}\times\dot{\mathbf{u}}]\quad\text{ on }S$ (18)
@@
[…]
Together (11), (16), (17) and (19) describe a coupled problem to be solved for the displacement $\mathbf{u}$ inside the grains and the normal stresses $p$ on the interfaces between grains.
@@ Appendix B Laplace transform
$\tilde{f}(s)=\int_{0}^{\infty}f(t)\mathrm{e}^{-st}\;\mathrm{d}t$ (20)
@@ Appendix C Weak forms
For finite element analysis the governing equations are most usefully written in weak form. Consider a domain $V$ potentially containing many individual grains. Taking the dot product of the complex conjugate of (11) with a vector test function $\mathbf{v}$, integrating over the domain $V$ and using Green’s identity yields
$\int_{V}\underline{\underline{\boldsymbol{\sigma}}}^{\mathbf{u}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}\;\mathrm{d}V-\int_{S_{i}}p^{*}[\mathbf{v}\cdot{\mathbf{n}}]\;\mathrm{d}S=\int_{S_{0}}{\mathbf{t}}^{0*}\cdot\mathbf{v}\;\mathrm{d}S$ (21)
where $S_{i}$ represents grain boundary surfaces; $S_{0}$ represents any exterior surfaces; and ${\mathbf{t}}^{0}$ any imposed tractions on those exterior surfaces. $\underline{\underline{\boldsymbol{\sigma}}}^{\mathbf{u}}$ is the stress tensor formed from the displacement $\mathbf{u}$. $\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}$ is the symmetric strain tensor formed from the test function $\mathbf{v}$. The colon (::::) denotes the tensor contraction on two indices. The terms involving the normal and tangential components of the traction at grain boundaries have been rewritten using (16) and (19). Superscript denotes complex conjugation. Note that both the displacement and the test function are complex functions, and the Laplace transform parameter $s$ may also be complex.
Multiplying the complex conjugate of (17) with a scalar test function $q$, integrating over the grain boundary surfaces and using Green’s identity yields
$-\int_{S_{i}}[\mathbf{u}^{*}\cdot{\mathbf{n}}]q\;\mathrm{d}S+\frac{\Omega\delta D^{\text{gb}}}{kTs^{*}}\int_{S_{i}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}q\;\mathrm{d}S=-\frac{\Omega\delta}{s^{*}}\int_{C}{\mathbf{j}_{\text{v}}^{*0}}\cdot\boldsymbol{\nu}\,q\;\mathrm{d}l$ (22)
@@ Appendix D Periodic homogenisation
$\mathbf{u}=\mathbf{U}+\hat{\mathbf{u}}$ (23)
where $\mathbf{U}$ represents a rigid body motion and $\hat{\mathbf{u}}$ represents the remaining elastic deformation. Each grain in the tessellation will be assumed to have the same elastic deformation $\hat{\mathbf{u}}$, but that $\mathbf{U}$ will vary in a systematic and prescribed manner. Following [8] (section 3), the kinematics of the rigid body motions will be prescribed in terms of macroscale strain measures $\underline{\underline{\boldsymbol{\Gamma}}}$ and $\underline{\underline{\boldsymbol{K}}}$, where the symmetric part of $\underline{\underline{\boldsymbol{\Gamma}}}$ represents the macroscale strain tensor. The jump in the rigid body displacement across the grain boundaries is [8, equation (21)]
$\left[\mathbf{U}\right]=\underline{\underline{\boldsymbol{\Gamma}}}\cdot\mathbf{R}+\left(\underline{\underline{\boldsymbol{K}}}\cdot\mathbf{R}\right)\times{\mathbf{d}}$ (24)
where $\mathbf{R}$ is a vector which joins the centroids of the neighbouring grains, and ${\mathbf{d}}$ is a vector joining the centroid of the neighbouring grain to the position on the grain boundary [8, Figure 1]. In the present work we will only consider upscaling to a Cauchy continuum, not a more general micropolar continuum as was done in [8]. We will thus neglect the strain measure $\underline{\underline{\boldsymbol{K}}}$ associated with gradients in microrotation, prescribing
$\left[\mathbf{U}\cdot{\mathbf{n}}\right]={\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}\cdot\mathbf{R},$ (25)
and solve for a vector displacement field $\hat{\mathbf{u}}$ which is identical in each grain in the periodic tessellation. Substitution of (23) and (25) into the weak forms (21) and (22) yields
$\int_{V}\underline{\underline{\boldsymbol{\sigma}}}^{\hat{\mathbf{u}}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}\;\mathrm{d}V-\int_{S_{i}}p^{*}[\mathbf{v}\cdot{\mathbf{n}}]\;\mathrm{d}S=0$ (26)
$-\int_{S_{i}}[\hat{\mathbf{u}}^{*}\cdot{\mathbf{n}}]q\;\mathrm{d}S+\frac{\Omega\delta D^{\text{gb}}}{kTs^{*}}\int_{S_{i}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}q\;\mathrm{d}S=\int_{S_{i}}\left({\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}^{*}\cdot\mathbf{R}\right)q\;\mathrm{d}S$ (27)
@@ D.1 Equivalent energetics
$\overline{\sigma}_{ij}=C_{ijkl}\Gamma_{kl}$ (28)
@@
$\Psi=\frac{1}{V}\left(\int_{V}\underline{\underline{\boldsymbol{\sigma}}}^{\mathbf{u}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{u}}\;\mathrm{d}V-\frac{\Omega\delta D^{\text{gb}}}{kTs^{*}}\int_{S_{i}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}p\;\mathrm{d}S\right)$ (29)
@@
$\Psi=\overline{\sigma}_{ij}\Gamma_{ij}=\Gamma_{ij}C_{ijkl}^{*}\Gamma_{kl}^{*}.$ (30)
@@ Appendix E Scaling
$\tau=\frac{kTd^{3}}{\delta D^{\text{gb}}\Omega\mu}$ (31)
@@ Appendix F Numerical Implementation
The computational domain consists of a single grain. A finite element mesh of the domain is constructed (triangular in 2D, tetrahedral in 3D). The variable $\hat{\mathbf{u}}$ is solved for throughout the domain, the variable $p$ solved for only on the exterior boundary of the domain. Periodic boundary conditions on $p$ are applied to impose invariance under translations by lattice vectors, where a point on each grain boundary is identical under translation to a point on the opposite face. Since the grain boundaries are on the edges of the computational domain, each section of domain boundary contains half the total grain boundary, but by periodicity the same boundary is traversed twice when integrating round the total surface of the grain. This simplifies the calculation of quantities involving jumps across the interface, as a single integral on a grain boundary involving a jump can be split into two separate integrals only involving displacements on one side of the boundary. The weak forms (32) and (33) can then be written as
$\int_{V_{\text{grain}}}\underline{\underline{\boldsymbol{\sigma}}}^{\hat{\mathbf{u}}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}\;\mathrm{d}V-\int_{S_{\text{grain}}}p^{*}\mathbf{v}\cdot{\mathbf{n}}\;\mathrm{d}S$ $=0$ (32)
$-\int_{S_{\text{grain}}}\hat{\mathbf{u}}^{*}\cdot{\mathbf{n}}\,q\;\mathrm{d}S+\frac{\Omega\delta D^{\text{gb}}}{2kTs^{*}}\int_{S_{\text{grain}}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}q\;\mathrm{d}S$ $=\frac{1}{2}\int_{S_{\text{grain}}}\left({\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}^{*}\cdot\mathbf{R}\right)q\;\mathrm{d}S$ (33)
@@ Appendix G Viscoelastic response functions
The main outcome of the finite element modelling is the tensor $C_{ijkl}$ which relates macroscale stress and strain in the Laplace-transformed variables of (28). For the simplest cases, this response is isotropic and can be written
$C_{ijkl}=s\tilde{K}(s)\delta_{ij}\delta_{kl}+s\tilde{G}(s)\left(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-\frac{2}{N}\delta_{ij}\delta_{kl}\right)$ (34)
@@
$s\tilde{J}(s)=\frac{1}{s\tilde{G}(s)}.$ (35)
@@
$G^{*}(\omega)=\left.s\tilde{G}(s)\right|_{s=i\omega}.$ (36)
@@
$J^{*}(\omega)=\left.s\tilde{J}(s)\right|_{s=i\omega}$ (37)
@@
$G^{*}(\omega)$ $=G_{1}(\omega)+iG_{2}(\omega),$ (38)
$J^{*}(\omega)$ $=J_{1}(\omega)-iJ_{2}(\omega).$ (39)
@@
$Q^{-1}(\omega)=\tan\arg G^{*}(\omega)=\frac{G_{2}(\omega)}{G_{1}(\omega)}=\frac{J_{2}(\omega)}{J_{1}(\omega)}.$ (40)
@@ H.1 Maxwell
$s\tilde{J}(s)=J_{0}\left(1+\frac{1}{s\tau_{M}}\right)$ (41)
@@
$J(t)=J_{0}\left(1+\frac{t}{\tau_{M}}\right).$ (42)
@@ H.2 Kelvin-Voigt
$s\tilde{J}(s)$ $=\frac{J_{r}}{1+s\tau_{r}},$ (43)
$J(t)$ $=J_{r}\left(1-\mathrm{e}^{-t/\tau_{r}}\right),$ (44)
@@ H.3 Burgers
$s\tilde{J}(s)$ $=J_{0}\left({1+\frac{\Delta}{1+s\tau_{r}}+\frac{1}{s\tau_{M}}}\right),$ (45)
$J(t)$ $=J_{0}\left(1+\Delta\left(1-\mathrm{e}^{-t/\tau_{r}}\right)+\frac{t}{\tau_{M}}\right),$ (46)

### v2

@@ A.1 Elasticity
$\nabla\cdot\underline{\underline{\boldsymbol{\sigma}}}=\mathbf{0}$ (A 11)
@@
$\sigma_{ij}=\lambda e_{kk}\delta_{ij}+2\mu e_{ij}$ (A 12)
@@
$e_{ij}=\frac{1}{2}\left(\frac{\partial u_{i}}{\partial x_{j}}+\frac{\partial u_{j}}{\partial x_{i}}\right).$ (A 13)
@@ A.2 Grain boundary diffusion
On the boundaries between grains the elastic displacement is discontinuous. Diffusion of atoms and vacancies along the grain boundaries leads to plating out or removal of material. Grain boundary diffusion is described by [e.g. 8, section 2.2]
$\left[\dot{\mathbf{u}}\cdot{\mathbf{n}}\right]+\Omega\delta D^{\text{gb}}_{\text{v}}\nabla_{\perp}^{2}c=0$ (A 14)
@@
$c=c_{0}\left(1+\frac{\Omega p}{kT}\right)$ (A 15)
@@
$p\equiv{\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\sigma}}}\cdot{\mathbf{n}}\quad\text{on }S,$ (A 16)
where $S$ is the grain-boundary interface. (A 15) can be substituted into (A 14) to write
$\left[\dot{\mathbf{u}}\cdot{\mathbf{n}}\right]+\frac{\Omega\delta D^{\text{gb}}}{kT}\nabla_{\perp}^{2}p=0,\\$ (A 17)
@@ A.3 Grain boundary sliding
${\mathbf{n}}\times\underline{\underline{\boldsymbol{\sigma}}}\cdot{\mathbf{n}}=\frac{\eta_{\text{gb}}}{\delta}[{\mathbf{n}}\times\dot{\mathbf{u}}]\quad\text{ on }S$ (A 18)
@@
${\mathbf{n}}\times\underline{\underline{\boldsymbol{\sigma}}}\cdot{\mathbf{n}}=\mathbf{0}\quad\text{on }S.$ (A 19)
Together (A 11), (A 16), (A 17) and (A 19) describe a coupled problem to be solved for the displacement $\mathbf{u}$ inside the grains and the normal stresses $p$ on the interfaces between grains.
@@ Appendix B Laplace transform
$\tilde{f}(s)=\int_{0}^{\infty}f(t)\mathrm{e}^{-st}\;\mathrm{d}t$ (B 20)
@@ Appendix C Weak forms
For finite element analysis the governing equations are most usefully written in weak form. Consider a domain $V$ potentially containing many individual grains. Taking the dot product of the complex conjugate of (A 11) with a vector test function $\mathbf{v}$, integrating over the domain $V$ and using Green’s identity yields
$\int_{V}\underline{\underline{\boldsymbol{\sigma}}}^{\mathbf{u}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}\;\mathrm{d}V-\int_{S_{i}}p^{*}[\mathbf{v}\cdot{\mathbf{n}}]\;\mathrm{d}S=\int_{S_{0}}{\mathbf{t}}^{0*}\cdot\mathbf{v}\;\mathrm{d}S$ (C 21)
where $S_{i}$ represents grain boundary surfaces; $S_{0}$ represents any exterior surfaces; and ${\mathbf{t}}^{0}$ any imposed tractions on those exterior surfaces. $\underline{\underline{\boldsymbol{\sigma}}}^{\mathbf{u}}$ is the stress tensor formed from the displacement $\mathbf{u}$. $\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}$ is the symmetric strain tensor formed from the test function $\mathbf{v}$. The colon (::) denotes the tensor contraction on two indices. The terms involving the normal and tangential components of the traction at grain boundaries have been rewritten using (A 16) and (A 19). Superscript denotes complex conjugation. Note that both the displacement and the test function are complex functions, and the Laplace transform parameter $s$ may also be complex.
Multiplying the complex conjugate of (A 17) with a scalar test function $q$, integrating over the grain boundary surfaces and using Green’s identity yields
$\int_{S_{i}}[\mathbf{u}^{*}\cdot{\mathbf{n}}]q\;\mathrm{d}S+\frac{\Omega\delta D^{\text{gb}}}{kTs^{*}}\int_{S_{i}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}q\;\mathrm{d}S=-\frac{\Omega\delta}{s^{*}}\int_{C}{\mathbf{j}_{\text{v}}^{*0}}\cdot\boldsymbol{\nu}\,q\;\mathrm{d}l$ (C 22)
@@ Appendix D Periodic homogenisation
$\mathbf{u}=\mathbf{U}+\hat{\mathbf{u}}$ (D 23)
where $\mathbf{U}$ represents a rigid body motion and $\hat{\mathbf{u}}$ represents the remaining elastic deformation. Each grain in the tessellation will be assumed to have the same elastic deformation $\hat{\mathbf{u}}$, but that $\mathbf{U}$ will vary in a systematic and prescribed manner. Following [9] (section 3), the kinematics of the rigid body motions will be prescribed in terms of macroscale strain measures $\underline{\underline{\boldsymbol{\Gamma}}}$ and $\underline{\underline{\boldsymbol{K}}}$, where the symmetric part of $\underline{\underline{\boldsymbol{\Gamma}}}$ represents the macroscale strain tensor. The jump in the rigid body displacement across the grain boundaries is [9, equation (21)]
$\left[\mathbf{U}\right]=\underline{\underline{\boldsymbol{\Gamma}}}\cdot\mathbf{R}+\left(\underline{\underline{\boldsymbol{K}}}\cdot\mathbf{R}\right)\times{\mathbf{d}}$ (D 24)
where $\mathbf{R}$ is a vector which joins the centroids of the neighbouring grains, and $\mathbf{d}$} is a vector joining the centroid of the neighbouring grain to the position on the grain boundary [9, Figure 1]. In the present work we will only consider upscaling to a Cauchy continuum, not a more general micropolar continuum as was done in [9]. We will thus neglect the strain measure $\underline{\underline{\boldsymbol{K}}}$ associated with gradients in microrotation, prescribing
$\left[\mathbf{U}\cdot{\mathbf{n}}\right]={\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}\cdot\mathbf{R},$ (D 25)
and solve for a vector displacement field $\hat{\mathbf{u}}$ which is identical in each grain in the periodic tessellation. Substitution of (D 23) and (D 25) into the weak forms (C 21) and (C 22) yields
$\int_{V}\underline{\underline{\boldsymbol{\sigma}}}^{\hat{\mathbf{u}}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}\;\mathrm{d}V-\int_{S_{i}}p^{*}[\mathbf{v}\cdot{\mathbf{n}}]\;\mathrm{d}S=0$ (D 26)
$-\int_{S_{i}}[\hat{\mathbf{u}}^{*}\cdot{\mathbf{n}}]q\;\mathrm{d}S+\frac{\Omega\delta D^{\text{gb}}}{kTs^{*}}\int_{S_{i}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}q\;\mathrm{d}S=\int_{S_{i}}\left({\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}^{*}\cdot\mathbf{R}\right)q\;\mathrm{d}S$ (D 27)
@@ D.1 Equivalent energetics
$\overline{\sigma}_{ij}=C_{ijkl}\Gamma_{kl}$ (D 28)
@@
$\Psi=\frac{1}{V}\left(\int_{V}\underline{\underline{\boldsymbol{\sigma}}}^{\mathbf{u}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{u}}\;\mathrm{d}V-\frac{\Omega\delta D^{\text{gb}}}{kTs^{*}}\int_{S_{i}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}p\;\mathrm{d}S\right)$ (D 29)
@@
$\Psi=\overline{\sigma}^{*}_{ij}\Gamma_{ij}=\Gamma_{ij}C_{ijkl}^{*}\Gamma_{kl}^{*}.$ (D 30)
@@ Appendix E Scaling
$\tau=\frac{kTd^{3}}{\delta D^{\text{gb}}\Omega\mu}$ (E 31)
@@ Appendix F Numerical Implementation
The computational domain consists of a single grain. A finite element mesh of the domain is constructed (triangular in 2D, tetrahedral in 3D). The variable $\hat{\mathbf{u}}$ is solved for throughout the domain, the variable $p$ solved for only on the exterior boundary of the domain. Periodic boundary conditions on $p$ are applied to impose invariance under translations by lattice vectors, where a point on each grain boundary is identical under translation to a point on the opposite face. Since the grain boundaries are on the edges of the computational domain, each section of domain boundary contains half the total grain boundary, but by periodicity the same boundary is traversed twice when integrating round the total surface of the grain. This simplifies the calculation of quantities involving jumps across the interface, as a single integral on a grain boundary involving a jump can be split into two separate integrals only involving displacements on one side of the boundary. The weak forms (F 32) and (F 33) can then be written as
$\int_{V_{\text{grain}}}\underline{\underline{\boldsymbol{\sigma}}}^{\hat{\mathbf{u}}*}:\underline{\underline{\boldsymbol{e}}}^{\mathbf{v}}\;\mathrm{d}V-\int_{S_{\text{grain}}}p^{*}\mathbf{v}\cdot{\mathbf{n}}\;\mathrm{d}S$ $=0$ (F 32)
$-\int_{S_{\text{grain}}}\hat{\mathbf{u}}^{*}\cdot{\mathbf{n}}\,q\;\mathrm{d}S+\frac{\Omega\delta D^{\text{gb}}}{2kTs^{*}}\int_{S_{\text{grain}}}\nabla_{\perp}p^{*}\cdot\nabla_{\perp}q\;\mathrm{d}S$ $=\frac{1}{2}\int_{S_{\text{grain}}}\left({\mathbf{n}}\cdot\underline{\underline{\boldsymbol{\Gamma}}}^{*}\cdot\mathbf{R}\right)q\;\mathrm{d}S$ (F 33)
@@ Appendix G Viscoelastic response functions
The main outcome of the finite element modelling is the tensor $C_{ijkl}$ which relates macroscale stress and strain in the Laplace-transformed variables of (D 28). For the simplest cases, this response is isotropic and can be written
$C_{ijkl}=s\tilde{K}(s)\delta_{ij}\delta_{kl}+s\tilde{G}(s)\left(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}-\frac{2}{N}\delta_{ij}\delta_{kl}\right)$ (G 34)
@@
$s\tilde{J}(s)=\frac{1}{s\tilde{G}(s)}.$ (G 35)
@@
$G^{*}(\omega)=\left.s\tilde{G}(s)\right|_{s=i\omega}.$ (G 36)
@@
$J^{*}(\omega)=\left.s\tilde{J}(s)\right|_{s=i\omega}$ (G 37)
@@
$G^{*}(\omega)$ $=G_{1}(\omega)+iG_{2}(\omega),$ (G 38)
$J^{*}(\omega)$ $=J_{1}(\omega)-iJ_{2}(\omega).$ (G 39)
@@
$Q^{-1}(\omega)=\tan\arg G^{*}(\omega)=\frac{G_{2}(\omega)}{G_{1}(\omega)}=\frac{J_{2}(\omega)}{J_{1}(\omega)}.$ (G 40)
@@ H.1 Maxwell
$s\tilde{J}(s)=J_{0}\left(1+\frac{1}{s\tau_{M}}\right)$ (H 41)
@@
$J(t)=J_{0}\left(1+\frac{t}{\tau_{M}}\right).$ (H 42)
@@ H.2 Kelvin-Voigt
$s\tilde{J}(s)$ $=\frac{J_{r}}{1+s\tau_{r}},$ (H 43)
$J(t)$ $=J_{r}\left(1-\mathrm{e}^{-t/\tau_{r}}\right),$ (H 44)
@@ H.3 Burgers
$s\tilde{J}(s)$ $=J_{0}\left({1+\frac{\Delta}{1+s\tau_{r}}+\frac{1}{s\tau_{M}}}\right),$ (H 45)
$J(t)$ $=J_{0}\left(1+\Delta\left(1-\mathrm{e}^{-t/\tau_{r}}\right)+\frac{t}{\tau_{M}}\right),$ (H 46)